OpenStudy (anonymous):

What is the largest prime factor of 36^2-17^2? What is the largest prime factor of 36^2-17^2? @Mathematics

OpenStudy (anonymous):

53

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

a^2 - b^2

OpenStudy (anonymous):

a^2-b^2 = (a-b)*(a+b) 36^2-17^2 = (36-17) *(36+17) = 19 * 53